You purchase 100 shares for $60 a share ($6,000), and after a year the price rises to $70. Calculate the percentage return on your investment if you bought the stock on margin and the margin requirement was (ignore commissions, dividends, and interest expense):

  1. 25 percent. Round your answer to one decimal place.

      %

  2. 65 percent. Round your answer to one decimal place.

      %

  3. 75 percent. Round your answer to one decimal place.

      %
Answer #1

=number of shares*(selling price-purchase price)/(number of shares*purchase price*margin %)

1.
=(100*(70-60))/(100*60*25%)
=66.67%

2.
=(100*(70-60))/(100*60*65%)
=25.64%

3.
=(100*(70-60))/(100*60*75%)
=22.22%